OpenStudy (ksaimouli):

let f be a function such that f''(x)=6x+12

OpenStudy (ksaimouli):

find f(X) if the graph of f is tangent to the line 4x-y=5 at the point (0,-5)

OpenStudy (ksaimouli):

@phi

OpenStudy (phi):

Did you integrate it ?

OpenStudy (ksaimouli):

no

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):

should i integrate 2 times?

OpenStudy (anonymous):

Yes.

OpenStudy (phi):

First step is find f'(x) by integrating once remember to include the constant

OpenStudy (ksaimouli):

3x^2+12x+c

OpenStudy (phi):

2nd step is figure out the constant using the info given

OpenStudy (ksaimouli):

-5=c

OpenStudy (mayankdevnani):

take the antiderivative of the equation f'(x) = 3x^2 + 12x + c the tangent line given is 4x - y = 5, or y = 4x - 5 so f'(x) = 4 when x = 0 use this information to solve for c 4 = 3(0) + 12(0) + c c = 4 plug in c to get f'(x) = 3x^2 +12x + 4 now take the antiderivative of this equation to get f(x) = x^3 + 6x^2 + 4x + c plug in the point (0, -5) to solve for c -5 = 0 + 0 + 0 + c c= -5 f(x) = x^3 + 6x^2 + 4x - 5

OpenStudy (phi):

-5=c NO the first derivative is the slope. you have to match the slope of the given line

OpenStudy (ksaimouli):

4x-y=5 ?

OpenStudy (phi):

yes and in slope-intercept form y= 4x-5 slope is 4

OpenStudy (ksaimouli):

okay

OpenStudy (ksaimouli):

what to do next

OpenStudy (phi):

3x^2 +12x +c = slope slope=4 at point (0,-5)

OpenStudy (ksaimouli):

so c=4

OpenStudy (phi):

yes. now you can integrate again to find f(x) remember the constant

OpenStudy (ksaimouli):

\[x^3+6x^2+4x+C\]

OpenStudy (ksaimouli):

now how to find c

OpenStudy (phi):

yes but you know a point on the curve (0, -5)

OpenStudy (ksaimouli):

c=-5

OpenStudy (phi):

yes, see maya's post for the computation in one write-up

OpenStudy (ksaimouli):

thanks how to find average value of f(x) on the closed interval (-1,1)

OpenStudy (phi):

find the area under the curve f(x) the average f(x) is the height that gives the same area height*width = area under the curve so integrate between -1 and 1 and then divide by the width ( 2)

OpenStudy (phi):

Does that make sense?

OpenStudy (ksaimouli):

hmm but is their aother instead of graphing it

OpenStudy (phi):

graphing? I would integrate and divide by 2

OpenStudy (ksaimouli):

got u what we got before (f(x)) should be integrated on -1 to 1 and divide by 2

OpenStudy (ksaimouli):

right?

OpenStudy (phi):

yes

OpenStudy (ksaimouli):

thx